Question 1
a) Doug (resident of Australia) earns $190,000 from his employment as an IT specialist. He has
no allowable deductions.
b) Doug (resident of Australia) earns $32,500 from his employment as an IT specialist. He has
no allowable deductions.
c) XYZ Pty Limited (resident of Australia and not an SBE / BRE) has assessable income of $32,500 for
the tax year. It has no allowable deductions.
Required:
Calculate the net tax payable / refundable that would be assessed in relation to the situations
referred to above for the year ended 30 June 2019.
1
Expert's answer
2020-12-10T13:40:38-0500

а)expected income=190 000

From $ 180,001 or more – $54 097. Plus 45 cents for every $1 of more than $ 180 000

54097+(190 000-180 000)*0,45=54097+4500=58597

Medicare: 190 000*1.50=2850

LITO and LMITO 0

net tax payable=58597+2850=61447

b)expected income=32 500

From $ 18 201 to $ 37 ,000 – 19 cents for every $1 over $18 200

(32500-18200)*0.19=2717

Medicare: 32500*0=0

LMITO: 250

LITO:700

net tax payable=2717-250-700=1767

c)expected income=32 500

from $ 18 201 to $ 37,000 – 19 cents for every $1 over $18 200

(32500-18200)*0.19=2717

Medicare: 32500*0=0

LMITO: 250

LITO:700

net tax payable=2717-250-700=1767